Shares of American Express (AXP) are currently valued at $117.72 per share. If the required return is 10.6% and dividends are growing at 9%, what is the amount of the current dividend the firm is paying?
Answer #1

Required return=(D1/Current price)+Growth rate

0.106=(D1/117.72)+0.09

D1=(0.106-0.09)*117.72

=$1.88352

Hence current dividend=D1*Present value of discounting factor(rate%,time period)

=1.88352/1.09

=$1.728